please help me solve this problem. 4divided by 7/10
Answer by Alan3354(69443)

please help me solve this problem. 4divided by 7/10
------------
When dividing by a fraction, invert the fraction and multiply by it.
4 divided by 7/10 = 4 * (10/7)
= 40/7